The perimeter of rectangularposter is included 80in. The length is 10 in. greater than the width. Find the length.
Answer by Leaf W.(135)

1. Name variables--
width: x
length: "the length is 10 in. greater than the width": width + 10: x + 10
2. Perimeter of a rectangle: 2(length) + 2(width)
Plug in variables named in Step 1: 2(x + 10) + 2(x)
In the problem, it states that the perimeter is 80 in., so put that equal to 80: 2(x + 10) + 2x = 80
3. Simplify and solve for x.
Distribute: 2x + 20 + 2x = 80
Combine like terms: 4x + 20 = 60
Subtract 20 from both sides: 4x = 60
Divide both sides by 4: x = 15
Therefore, x (the width) is 15. However, the problem is looking for the length. To find the length, plug this value for x (15) into the expression for length (x + 10) found in Step 1.
15 + 10
25
Therefore, the length of the rectangular poster is 25 inches.